OpenStudy (anonymous):

Let f(x) = 2x - 6 Solve f-1(x) when x = 2.

jimthompson5910 (jim_thompson5910):

\[\large f(x) = 2x - 6\] \[\large y = 2x - 6\] \[\large x = 2y - 6\] \[\large x+6 = 2y\] \[\large \frac{x+6}{2} = y\] \[\large y = \frac{x+6}{2}\] \[\large f^{-1}(x) = \frac{x+6}{2}\] \[\large f^{-1}(2) = \frac{2+6}{2}\] \[\large f^{-1}(2) = ???\]

OpenStudy (anonymous):

So, the answer would be -2 ?

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

4?

jimthompson5910 (jim_thompson5910):

you need to evaluate \(\large \frac{2+6}{2}\)

OpenStudy (anonymous):

8 over 2 which equals 4?

OpenStudy (anonymous):

2 + 6 = 2 if you divide that by 2 that would be 6 ? but, thats not an answer choice @jim_thompson5910 sorry, im a bit confused ..

jimthompson5910 (jim_thompson5910):

\(\large \frac{2+6}{2} = \frac{8}{2} = 4\)

jimthompson5910 (jim_thompson5910):

not sure how you're getting 2+6 = 2

OpenStudy (anonymous):

thats what i said! 4..

jimthompson5910 (jim_thompson5910):

yes I know henry

OpenStudy (anonymous):

I suck at math, but thankyou !!

jimthompson5910 (jim_thompson5910):

I'm writing it for Nurialozza96

OpenStudy (anonymous):

lol i should get a Best Response huh

jimthompson5910 (jim_thompson5910):

I'm sure she's just fine at it, she just needs more practice, that's all

OpenStudy (anonymous):

@henryarias5 you're childish, get off my page please. goodbye. and @jim_thompson5910 thankyouuuu !!!

jimthompson5910 (jim_thompson5910):

you're welcome